OpenStudy (anonymous):

A 50-foot tree casts a shadow 9 feet long. The sine of the angle of elevation of the top of the tree to the sun is approximately _____. A. .98 B. .01 C. .18 D. .02

OpenStudy (anonymous):

 \[h^2=50^2+9^2\] \[h=\sqrt{50^2+9^2}\] \[h=50.80354318\] \[\sin (A) = 9/h\] \[\sin(A)=9/50.80354318=.1771529983=.18\] Ans: C.

OpenStudy (anonymous):

The answer was .98! But thanks for your help!